# Connect the Dots: Knowledge Graph–Guided Crawler Attack on Retrieval-Augmented Generation Systems

### v1

@@ Abstract
Retrieval-augmented generation (RAG) systems integrate document retrieval with large language models and have been widely adopted. However, in privacy-related scenarios (e.g., medical diagnosis), RAG introduces a new privacy risk: adversaries can issue carefully crafted queries to exfiltrate sensitive content from the underlying corpus gradually. Although recent studies have demonstrated multi-turn extraction attacks, they rely on heuristics and fail to perform long-term extraction planning. As a result, they typically stagnate after a few rounds and exfiltrate only a small fraction of the private corpus. To address these limitations, we formulate the RAG extraction attack as an adaptive stochastic coverage problem (ASCP). In ASCP, each query is treated as a probabilistic action that aims to maximize conditional marginal gain (CMG), enabling principled long-term planning under uncertainty. However, integrating ASCP with practical RAG attack faces three key challenges: unobservable CMG, intractability in the action space, and feasibility constraints. To overcome these challenges, we maintain a global attacker-side state to guide the attack. Building on this idea, we introduce RAGCrawler, which builds a knowledge graph to represent revealed information, uses this global state to estimate CMG, and plans queries in semantic space that target unretrieved regions. It also generates benign-looking prompts that elicit the remaining content while preserving stealth. In comprehensive experiments across diverse RAG architectures and datasets, our proposed method, RAGCrawler, consistently outperforms all baselines. It achieves up to 84.4% corpus coverage within a fixed query budget and deliver an average improvement of 20.7% over the top-performing baseline. It also maintains high semantic fidelity and strong content reconstruction accuracy with low attack cost. Crucially, RAGCrawler proves its robustness by maintaining effectiveness against advanced RAG systems employing query rewriting and multi-query retrieval strategies. Our work reveals significant security gaps and highlights the pressing need for stronger safeguards for RAG.
@@ 1 Introduction
Retrieval-Augmented Generation (RAG) has emerged as a powerful paradigm that integrates information retrieval with large language models (LLMs), allowing systems to dynamically access and utilize external knowledge bases for tasks such as question answering, summarization, and decision support [karpukhin2020dense, lewis2020retrieval]. By providing timely, domain-specific evidence beyond a model’s parametric memory, RAG has seen rapid adoption across settings ranging from enterprise knowledge management [xu2024generative] to healthcare [al2023transforming], and customer service [loukas2023making].
RAG’s retrieval capabilities introduce significant privacy risks, as many of the data sources it accesses or stores contain confidential information. Recent studies demonstrate that sensitive data can leak from RAG by eliciting verbatim passages [zeng2024good, qifollow], exposing personal identifiers from clinical notes or confidential clauses from internal documents. Such leaks can lead to non-compliance with GDPR and HIPAA [voigt2017eu, annas2003hipaa], resulting in legal, financial, and reputational harm [lasso2024ragsecurity]. Abusing the retrieval capabilities, adversaries can exploit the retriever as a high-precision oracle to exfiltrate private corpus content through carefully crafted prompts.
Understanding of the threat posed by extraction attacks on RAG remains limited, as existing studies often overlook adversaries’ global planning capabilities when crafting the prompts. In known attacks, prompts often build on recent responses, expanding on the retrieved context to uncover progressively more information. For example, an attack might extract a keyword from a recent answer and query it further [wang2025silent], or instruct the LLM to elaborate on a just-mentioned fact [jiang2024rag]. These heuristic methods lack global planning and will often drifts off topic (as shown in the left panel Fig. 1(a)) or re-touches previously seen content (as shown in the middle panel Fig. 1(a)), producing low-value responses and stalling early, leaving a substantial portion of the corpus remains unexplored as illustrated in Fig 1(b). Most importantly, while the empirical results demonstrate the feasibility of extracting content from RAG, the theoretical grounding and provable guarantees remain missing.
In this work, we provide theoretical guarantees for data extraction and systematical approach to analysis the extraction capability of practical adversaries. Our key insight is that these extraction attacks can be formalized as an instance of the Adaptive Stochastic Coverage Problem (ASCP) [nemhauser1978analysis, wolsey1982analysis, golovin2011adaptive], where different attacks can be viewed as various strategies for solving the underlying coverage objective. ASCP is well studied in submodular optimization and the near-optimal strategy exists: if the coverage function, representing the fraction of the corpus revealed, is both adaptively monotone and adaptively submodular, then a greedy policy that, at each step, selects the query with the largest conditional expected marginal gain (CMG) is a near-optimal strategy. In our formulation, each query is treated as an action that stochastically reveals a subset of hidden documents, and the objective is to maximize the expected corpus coverage given a limited query budget. Leveraging results from ASCP, we show that an attack on RAG, which globally plans queries to maximize CMG, can approach the theoretical upper bound on extracting RAG’s content.
Translating theoretical understanding into real-world attacks is challenging because of the following gaps that arise when instantiating the conditional margin gain (CMG), action space, and budget constraints of ASCP for RAG attacks: ❶ CMG of the attacker’s action is unobservable: the attacker cannot directly measure the coverage increment produced by a query and subsequent retrievals, since they lack a global view of the document corpus. Consequently, the true CMG is unobservable and cannot be computed exactly. ❷ Attacker action spaces are intractable: the query space $Q$ is effectively unbounded (any natural-language string). Exhaustive search for the CMG-maximizing query is infeasible absent additional structure or strong priors. ❸ Defense of real-word RAG poses its own feasibility constraints: queries in real-world RAG must remain natural and innocuous to avoid detection or refusal; high-gain intents must be expressed in policy-compliant surface forms that blend into ordinary usage. In addition, as real-world RAG systems often employ advanced query processing techniques such as query rewriting and multi-query retrieval, it becomes even more critical to articulate the intent explicitly to prevent the original meaning from being altered during these processes.
We propose a novel attack framework, RAGCrawler, to address these challenges. RAGCrawler maintains a structured attacker-side knowledge graph (KG) to give the attacker an overview of acquired information. With this global view, our attack executes a three-stage loop. First, KG construction utilizes existing answers and a coverage proxy to map graph growth to an estimate of coverage marginal gain, making it observable. Second, strategy scheduling replaces the intractable search with a two-stage selection process, driven by a score that fuses historical gain with structural exploration. Third, the query generation produces top-ranked anchors into natural, policy-compliant queries with history-aware deduplication. These three steps together act as an adaptive greedy coverage maximization strategy for RAGCrawler, which guarantees that RAGCrawler achieves near-optimal coverage of the RAG content.
To evaluate the effectiveness of RAGCrawler, we conduct a comprehensive evaluation across four benchmarks, four RAG generators (with safeguard variants), and two RAG retrievers. Results show that RAGCrawler consistently outperforms baselines in coverage rate, semantic fidelity, and reconstruction fidelity. RAGCrawler achieves an average coverage rate of 66.80%, which is a 20.70% improvement over the best baseline. Besides, the extracted content is of high quality, enabling the construction of a substitute RAG system that achieves a 0.6992 semantic similarity to the original RAG’s responses. RAGCrawler maintains effective against generators with safeguards (reaching an average coverage rate of 55.30%) and potential defense mechanisms (including query rewriting and multi-query retrieval). In summary, our contributions are:
  • We formalize data extraction attack in RAG as RAG Crawling, an adaptive stochastic coverage problem, and provide a rigorous theoretical foundation for the attack.
@@
  • We design RAGCrawler, a novel coverage-guided crawler that offers the provable near-optimal attack strategy under real-world constraints.
@@
  • We conduct extensive evaluations across diverse settings to demonstrate the effectiveness of RAGCrawler, as well as against modern RAG defenses.
@@
Figure 1: A motivating example. (a) A data extraction scenario on a private medical corpus. The continuation strategy follows the recent answer’s context, which can drift away from the corpus content; the keyword strategy reuses extracted keywords to form new queries, often yielding redundant information from the same semantic region; our proposed strategy maintains a global graph, detects unrevealed facts and targets them with new queries. (b) Knowledge coverage vs. query number for each strategy on the example corpus. (c) Distribution of retrieved facts in the corpus’ semantic space. Each point is a fact or document, with colors indicating queries from different strategies.
2 A Motivating Example
Consider an attacker targeting a private medical corpus, as shown in Fig. 1(a). The corpus contains de-identified records (represented by [Uncaptioned image], [Uncaptioned image], and [Uncaptioned image]). The attacker engages the RAG system in a dialogue and gradually coaxes out sensitive medical facts about each individual via carefully crafted follow-up queries. Previous results already show that even after removing direct personal identifiers, inference attacks can still re-identify individuals and reveal sensitive attributes when sufficient information is obtained [narayanan2008robust, sweeney2000simple, cai2025can]. This scenario (though synthetic for illustration) reflects a real threat model: by asking seemingly innocuous questions in sequence, an attacker can piece together confidential information about different people from the underlying private database. As illustrated in Fig. 2, gaps in knowledge by identifying insufficiently covered areas can be detected by a knowledge graph and then used to generate targeted queries to fill those gaps, delving deeper into the sensitive medical facts related to each individual.
Figure 2: Illustration of gap detection by the Global Strategy (Ours). The strategy identifies from the graph that the patient Refer to caption has a missing Personal Medical History, and guides the next exploration step to complete the knowledge.
@@ 2 A Motivating Example
Existing work can be categorized into two types: continuation expansion and keyword expansion. In the first strategy (e.g., RAGThief [jiang2024rag]), the new query is built on the context of existing answers. The problem of this strategy is that the query can easily drift away from the actual corpus content. For example, in Fig. 1(a), the last answer reveals [Uncaptioned image]’s family medical history and mentions the baby’s grandfather [Uncaptioned image]. A continuation-based attacker will ask about the grandfather next, but since [Uncaptioned image] is not in the corpus, this line of questioning yields no useful information. A keyword-based strategy (e.g., IKEA [wang2025silent]) pivots each new query around key terms gleaned from previous answers, but the problem is that it tends to stay within the same semantic neighborhood of the corpus. For example, in Fig. 1(a), keyword-based attack repeatedly ask about [Uncaptioned image]’s disease and does not explore new topics.
@@
Our work uses a growing knowledge graph to represent discovered entities and relations. A node represents an entity (e.g. a patient, a symptom), and edges capture relations or co-occurrences between entities (for example, patient $\rightarrow$ symptom if a patient is noted to have a certain symptom). As the attacker obtains new answers, the graph is expanded with any newly revealed entities or connections. By maintaining this global knowledge state, the attacker can detect gaps in what has been revealed so far. As illustrated in Fig. 2, the knowledge graph dynamically reveals incomplete regions of information. For instance, the graph shows that the patient [Uncaptioned image] has no extracted Personal Medical History, while the other patients [Uncaptioned image] and [Uncaptioned image] already have their medical histories identified. Then the next query will be directed to the uncovered data of [Uncaptioned image]’s unobserved medical history, and expanding overall coverage.
Fig. 1(b) shows the knowledge coverage rate v.s. the number of queries for our attack and two baseline strategies. Guided by the global knowledge graph, our attack (green line) achieves a much steeper increase in coverage compared to the continuation-based (orange curve; focusing on unrelated queries) and keyword-based (blue curve; focusing on repetitive queries) strategies.
@@
Fig. 1(c) provides a qualitative view of how each strategy explores the semantic space of the corpus. In this visualization, each point represents a distinct fact or document. The color represents the strategy that retrieves the document. The black box represents the space of the private corpus. We can observe that, guided by the global knowledge graph, our attack spread widely across the private space and covers many different clusters of information. The continuation-based queries (orange points) follow a narrow path through the space, gradually drifting away from the space center. The keyword-based queries (blue points) form a few dense clusters around specific anchor topics.
@@ 3 Background
In this section, we introduce the background, including Retrieval-Augmented Generation Systems (Sec. 3.1), Adaptive Stochastic Coverage Problem (Sec. 3.2) and Knowledge Graph (Sec. 3.3).
@@ 3.1 Retrieval-Augmented Generation Systems
where $sim(\cdot)$ represents the similarity function, and $\arg\text{ top-k}$ selects the top-k documents with the highest similarity scores. The generator $\mathcal{G}$ then generates an output based on the contextual information from the retrieved documentsThe generator then produces an answer conditioned on the query and retrieved context [lewis2020retrieval]:
@@
where $\text{wrapper}(\cdot,\cdot)$ denotes the system prompt that interleaves $q$ with the retrieved evidence.
@@
Query rewriting [lin2020conversational, ma2023query, mo2023convgqr, wang2025maferw] transforms the original query to improve retrievability, resolve ambiguity, repair typos, or strip unsafe instructions. The downstream retrieval and generation processes then operate on the rewrited query.
Multi-query retrieval [multiquery, li2024dmqr] generates multiple paraphrases of the original query and performs independent retrieval for each. The retrieved results are then aggregated into a candidate pool. Instead of concatenating all retrieved documents, RAG systems typically re-rank them to select the most relevant ones, with Reciprocal Rank Fusion (RRF) [cormack2009reciprocal] being a widely adopted method. Multi-query retrieval is not a dedicated defense, but its altered retrieval dynamics may affect the attack surface.

### v2

@@ Abstract
Stealing attacks pose a persistent threat to the intellectual property of deployed machine-learning systems. Retrieval-augmented generation (RAG) intensifies this risk by extending the attack surface beyond model weights to knowledge base that often contains IP-bearing assets such as proprietary runbooks, curated domain collections, or licensed documents. Recent work shows that multi-turn questioning can gradually steal corpus content from RAG systems, yet existing attacks are largely heuristic and often plateau early. We address this gap by formulating RAG knowledge-base stealing as an adaptive stochastic coverage problem (ASCP), where each query is a stochastic action and the attacker’s goal is to maximize the conditional expected marginal gain (CMG) in corpus coverage under a query budget. Bridging ASCP to real-world black-box RAG knowledge-base stealing raises three challenges: CMG is unobservable, the natural-language action space is intractably large, and feasibility constraints require stealthy queries that remain effective under diverse architectures. We introduce RAGCrawler, a knowledge graph-guided attacker that maintains a global attacker-side state to estimate coverage gains, schedule high-value semantic anchors, and generate non-redundant natural queries. Across four corpora and four generators with BGE retriever, RAGCrawler achieves 66.8% average coverage (up to 84.4%) within 1,000 queries, improving coverage by 44.90% relative to the strongest baseline. It also reduces the queries needed to reach 70% coverage by at least $4.03\times$ on average and enables surrogate reconstruction with answer similarity up to 0.699. Our attack is also scalable to retriever switching and newer RAG techniques like query rewriting and multi-query retrieval. These results highlight urgent needs to protect RAG knowledge assets.
@@ 1 Introduction
Intellectual property (IP) theft remains a central concern across the AI service landscape. A large body of work on model-stealing attacks and defenses [juuti2019prada, oliynyk2023know, carlini2024stealing, rakin2022deepsteal, orekondy2019knockoff, shen2022model, tramer2016stealing, he2021stealing, liu2022stolenencoder, yuan2022attack, das2025siamese, danda2021towards, orekondy2020prediction, li2022defending, kariyappa2020defending, wu2024efficient] underscores the urgency of protecting proprietary AI assets and preventing information theft. As modern AI service architectures grow increasingly complex, the IP attack surface extends beyond the model itself, requiring immediate investigation and comprehensive defense strategies.
Retrieval-Augmented Generation (RAG), now a ubiquitous architecture grounded in proprietary data, extends the attack surface to the retrieval corpus that sits on the path of every response. RAG couples a generator with a retriever over a document collection, using retrieved documents to improve factuality and domain adaptation [karpukhin2020dense, lewis2020retrieval]. In enterprise and vertical deployments, the corpus often contains IP-bearing assets (e.g., proprietary runbooks and policies, curated domain data, and licensed materials) whose acquisition and maintenance are costly [xu2024generative, al2023transforming, loukas2023making]. Because each response is grounded in retrieved documents, the public chat or API interface [vertexaisearch_pricing] becomes an attack surface: an attacker can progressively steal more of the underlying corpus using curated queries. Stealing a sufficiently large portion of the corpus can be enough to replicate the service: an attacker can pair the stolen corpus with an off-the-shelf LLM to build a substitute RAG assistant that approximates the victim’s behavior without access to the original deployment [wang2025silent].
Recent studies highlight that knowledge-base stealing poses an important and growing threat to deployed RAG systems, with direct copyright, licensing, and privacy implications for IP-bearing corpora [zeng2024good, qifollow, jiang2024rag, wang2025silent]. Attackers can elicit verbatim or near-verbatim spans from RAG outputs [zeng2024good, qifollow], and sustained multi-turn interaction can cumulatively expand the portion of a hidden corpus that becomes exposed [wang2025silent, jiang2024rag]. However, existing attacks typically generate prompts by locally reacting to the most recent response, either continuing the thread or pivoting on extracted keywords [jiang2024rag, wang2025silent]. Such heuristics often drift off-topic (left panel of Fig. 1 (a)) or revisit already-exposed regions (middle panel of Fig. 1 (a)), wasting query budget and plateauing early, leaving much of the corpus unexplored (Fig. 1 (b)). Moreover, while prior empirical studies demonstrate the knowledge-base theft, theoretical grounding and provable guarantees for coverage-maximizing attacks remain unknown. This motivates a critical question: can the adversary steal the data more efficiently, and where is the limit?
To address this question, we identify the fundamental gap in existing work: lack of a principled objective to maximize global corpus coverage. Without such an objective, query selection is driven by local reactions to the latest response, providing no globally consistent criterion to avoid redundancy and systematically expand coverage across turns. An effective knowledge-base stealing strategy should instead exploit the global state accumulated over the entire interaction history and choose each query by its expected marginal contribution to previously unseen corpus content.
We formalize this global objective by casting RAG knowledge-base stealing as RAG Crawling, an instance of the Adaptive Stochastic Coverage Problem (ASCP) [nemhauser1978analysis, wolsey1982analysis, golovin2011adaptive]. In our reduction, each query is an action that stochastically reveals a subset of hidden documents through retrieval, and the attacker aims to maximize the expected coverage of unique corpus items under a fixed query budget. ASCP provides an adaptive-greedy benchmark that selects the query with the largest conditional expected marginal gain (CMG) at each step. We further prove that RAG Crawling satisfies adaptive monotonicity and adaptive submodularity; thus, under the same budget of $B$ queries, the CMG-based adaptive greedy attack is provably near-optimal, achieving a (11/e)(1-1/e)-approximation to the expected corpus coverage attainable by an optimal adaptive attacker.
However, implementing this theory into a practical black-box attack faces three key instantiation challenges. ❶ CMG is unobservable: the attacker cannot directly measure the coverage increment induced by a query and its retrieval outcome without corpus access, so the true CMG cannot be exactly computed. ❷ The action space is intractable: the query space $Q$ is effectively unbounded (any natural-language string), making it infeasible to exhaustively search for the CMG-maximizing query without additional structure. ❸ Real-world feasibility constraints are strict: queries must remain natural and innocuous to avoid detection or refusal, and high-gain intents must be expressed in policy-compliant surface forms that remain stable under query rewriting and multi-query retrieval.
To solve these challenges, we introduce RAGCrawler, a novel attack framework consisting of three stages: knowledge graph construction, strategy scheduling, and query generation, that together overcome the above challenges. RAGCrawler maintains an attacker-side knowledge graph (KG) that provides a global summary of acquired information and supports planning. With this global state, RAGCrawler first uses KG growth in KG construction phase to estimate coverage gain and CMG. RAGCrawler next prunes the action space by selecting high-value semantic anchors in strategy scheduling phase. These semantic anchors are determined based on historical gains and the structural exploration of the KG. During the query generation phase, RAGCrawler enforces real-world feasibility by turning top-ranked anchors into fluent, policy-compliant natural-language queries with history-aware deduplication. Together, these three stages form a practical instantiation of the adaptive greedy policy for ASCP.
Experiments across diverse deployments show that RAGCrawler substantially improves RAG knowledge-base stealing effectiveness. Across four datasets and four RAG generators (including safeguard variants), it achieves up to 84.4% corpus coverage within a 1,000-query budget with BGE retriever. It improves coverage by 44.90% relative to the best baseline and reduces the queries needed to reach 70% coverage by at least 4.03× on average. The stolen content also supports service replication: a surrogate RAG built from the recovered corpus reaches up to 0.699 answer similarity to the victim system. RAGCrawler is scalable to other RAG techniques like retriever switching, query rewriting, and multi-query retrieval.
Our contribution.
  • Theory. We formalize RAG knowledge-base stealing as RAG Crawling, an adaptive stochastic coverage problem. We provide a principled, objective, and theoretical grounding for coverage-maximizing black-box attacks.
@@
  • System. We design RAGCrawler, a knowledge graph–guided crawler that instantiates the adaptive greedy coverage principle under real-world constraints.
@@
  • Evaluation. We conduct extensive evaluations across diverse settings to demonstrate the effectiveness of RAGCrawler, as well as against modern RAG defenses.
2 Motivating Example
Consider an attacker targeting a paid support troubleshooting RAG assistant, which is backed by a vendor support portal (Fig. 1(a)). Such portals are often built on private issue tracking tickets and internal resolution notes, and are usually considered IP assets [jira]. In our example, we will select three representative issues and assign an icon to each. (storage), (replication) , and (authentication) represent three issues. The attacker’s goal is to steal as much of the hidden knowledge base as possible through multi-round queries. By repackaging the stolen knowledge, attackers can mount further downstream attacks such as rehosting the RAG system, bypassing access controls, undercutting subscriptions, or violating licensing and trade secret protections [wang2025silent].
@@ 2 Motivating Example
Figure 1: A motivating example. (a) A RAG knowledge-base stealing scenario on an IP-bearing troubleshooting corpus. Upper answers represent knowledge obtained from earlier attack rounds. The continuation strategy follows the recent answer, which can drift away from the corpus; the keyword strategy pivots on extracted keywords, often yielding redundant information from the same semantic region; our proposed strategy maintains a global graph, detects unrevealed facts and targets them with new queries. (b) Knowledge coverage w.r.t query number for each strategy on an example corpus. (c) Distribution of retrieved document in the corpus’ semantic space. Each point is a document, with colors indicating queries from different strategies.
@@
Existing stealing attacks mainly rely on two local heuristics, continuation expansion and keyword expansion. In the first strategy (e.g., RAGThief [jiang2024rag]), the new query is built on the context of existing answers. The problem of this strategy is that the query can easily drift away from the actual corpus content. For example, in Fig. 1 (a), the last answer about Issue  attributes the failure to a dependency package . A continuation-based attacker will ask about next, but since it is not in the licensed knowledge base’s scope, this line of questioning yields no useful information. A keyword-based strategy (e.g., IKEA [wang2025silent]) pivots each new query around key terms of previous answers, keeping it mostly within the same semantic neighborhood of the corpus. For example, in Fig. 1 (a), the keyword-based attack repeatedly asks about ’s root cause and does not explore new contents.
Figure 2: Illustration of gap detection by the Global Strategy (Ours). The strategy identifies from the graph that the storage issue ( ) has a missing Hotfix facet, and guides the next exploration step to complete this unrevealed knowledge.
@@
Our approach maintains a growing knowledge graph that summarizes discovered entities and relations for globally-planned exploration. Fig. 2 shows the knowledge graph of our attack. Nodes represent entities (e.g., issues, hotfixes, modules), and edges capture relations or co-occurrences (e.g., issue $\rightarrow$ hotfix). As new answers arrive, the graph is expanded with all new entities or relations. By maintaining this global knowledge state, the attacker can detect gaps in what has been revealed so far. As illustrated in Fig. 2, the knowledge graph dynamically reveals incomplete regions of information. For instance, the graph shows that the issue has no revealed Hotfix, while the other issues ( and ) already have their hotfix information identified. Then the next query will be directed to the uncovered data of ’s unobserved hotfix steps, expanding overall coverage.
@@
Our empirical results show that using this global knowledge graph significantly improves the coverage rate. Fig. 1 (b) shows the knowledge coverage rate vs. the number of queries for our attack and two baseline strategies. Guided by the global knowledge graph, our attack (green line) converges much faster than continuation-based (orange curve; focusing on unrelated queries) and keyword-based (blue curve; focusing on repetitive queries) strategies. Fig. 1 (c) provides a qualitative view of how each strategy explores the semantic space of the corpus. Each point denotes a document, and the black box indicates the knowledge-base space. Guided by the global knowledge graph, our attack (green points) spread widely across the space and covers many different clusters of information. The continuation-based queries (orange points) follow a narrow path through the space, gradually drifting away from the space center. Keyword-based queries (blue points) form a few dense clusters around specific topics.
@@ 3.1 Retrieval-Augmented Generation Systems
where $sim(\cdot)$ represents the similarity function, and $\arg\text{ top-k}$ selects the top-k documents with the highest similarity scores. The generator $\mathcal{G}$ then produces an answer conditioned on the query and retrieved context [lewis2020retrieval]:
@@
where $\text{wrapper}(\cdot,\cdot)$ denotes the system prompt that interleaves $q$ with the retrieved documents.
@@
Query rewriting [lin2020conversational, ma2023query, mo2023convgqr, wang2025maferw] transforms the original query to improve retrievability, resolve ambiguity, repair typos, or strip unsafe instructions. The downstream retrieval and generation processes then operate on the rewritten query.
Multi-query retrieval [multiquery, li2024dmqr] generates multiple paraphrases of the original query and retrieves independently for each, aggregating the results into a candidate pool. Rather than concatenating all retrieved documents, RAG systems typically re-rank candidates to select the most relevant ones, with Reciprocal Rank Fusion (RRF) [cormack2009reciprocal] widely used. While not a dedicated defense, multi-query retrieval’s altered retrieval dynamics may change the attack surface.